# Inference on multiple quantiles in regression models by a rank-score approach

### v1

@@ Abstract
This paper tackles the challenge of performing multiple quantile regressions across different quantile levels and the associated problem of controlling the familywise error rate, an issue that is generally overlooked in practice. We propose a multivariate extension of the rank-score test and embed it within a closed-testing procedure to efficiently account for multiple testing. Theoretical foundations and simulation studies demonstrate that our method effectively controls the familywise error rate while achieving higher power than traditional corrections, such as Bonferroni.
@@ 1 Introduction
We focus on the problem of testing the evidence of a covariate effect across multiple quantile levels. In the literature, this is often referred to as regional quantile testing when the quantile levels form a continuous range, and as global or partial quantile regression when the levels form a finite set of quantiles $\{\tau_{1},\dots,\tau_{K}\}$. Various methods exist for regional quantile testing, with examples found in the work of sun2021model; belloni2019conditional; he2010detection. In contrast, a valid approach for conducting quantile regressions at multiple discrete $K$ levels is lacking. Typically, researchers perform $K$ separate regressions, analyzing each regression’s results and the corresponding $p$-values/test statistics independently.
@@
This practice, however, leads to a multiple testing problem (goeman2014multiple), where performing $K$ separate tests increases the risk of false discoveries if $p$-values are not adjusted to control family errors such as familywise error (FWER) or false discovery rates. The multiplicity problem is not commonly addressed within the quantile regression literature. To the best of our knowledge, only mrkvicka2023globalquantileregression consider multiple testing in this framework; yet, their proposal lacks a solid theoretical foundation.
@@
The strategy of applying quantile regression across various quantiles and making inferences about the effects of covariates without addressing multiple testing issues is a common oversight in various disciplines. Numerous examples can be found in studies spanning markets (wang2012computing), biology (agarwal2019quantile), management (bien2016study; karimi2024r), economics (fousekis2005demand; falk2012quantile), public health (juvanhol2016factors), health (cheah2022malaysian), education (bassett2002quantile), energy (cheng2020quantile), and climate (chamaille2007detecting). This widespread practice highlights a methodological gap, emphasizing the need for the field to develop and apply rigorous statistical methods that properly account for multiple comparisons and ensure the validity of the findings (ioannidis2005most).
@@
We propose a principled yet straightforward framework for performing valid inference across multiple quantile levels. A naive but valid approach would be to apply standard corrections, such as Bonferroni, to the set of $p$-values obtained from the $K$ separate quantile regressions. However, as demonstrated by goeman2021only, only closed testing procedures are admissible for strong FWER control; every other method is either equivalent to, or uniformly dominated by, some closed testing procedure. Consequently, we adopt the closed testing principle as the inferential backbone of our approach. The closed testing framework controls the FWER by testing all possible intersections of the individual hypotheses $H_{0}^{1},\dots,H_{0}^{K}$ using valid local tests. In short, for any $j=1,\dots,K$, we can reject $H_{0}^{j}$ at a pre-determined level $\alpha$ if all intersection hypotheses containing $H_{0}^{j}$ are rejected at that level $\alpha$. The validity and efficiency of the closed testing procedure thus hinge on the choice of these local tests. The commonly used Wald-type test for quantile regression (koenker1996quantile) suffers from finite-sample distortions, especially when heteroscedasticity affects the data, and slow convergence to its asymptotic distribution, even under independent and identically distributed errors (de2025closed; koenker2005quantile).
@@
A robust alternative is the rank-score test (gutenbrunner1993tests; koenker1999goodness; kocherginsky2005practical), whose multivariate version, however, has not yet been fully developed and theoretically investigated. We therefore propose the multivariate extension of the rank-score test, and we integrate it as the local test within the closed testing procedure. This approach increases analytical flexibility and guarantees valid control of type I errors across multiple quantile levels, thereby improving the reliability of statistical inferences from quantile regression analyses.
@@
To facilitate the application of the proposed methodology, we provide an R package named quasar (QUAntile Selective inference Adjustment via Rank-scores) available on CRAN (https://cran.r-project.org/web/packages/quasar/index.html).
@@
The manuscript is organized as follows. Section 2 presents the necessary background on quantile regression, while Section 3 introduces the closed testing procedure. Section 4 details the theory of the multivariate version of the rank-score test. Section 5 presents simulation studies showing that the Wald-type test fails to control the type I error rate, whereas the proposed method effectively controls the FWER while maintaining competitive power compared to traditional corrections such as Bonferroni.
@@ 2 Setting
We assume that the following location-shift model (koenker1999goodness) generates the data:
@@
where $\beta,\theta\in\mathbb{R}$, $\gamma,\eta\in\mathbb{R}^{p}$, and $U_{ni}$ are independent and identically distributed errors from a distribution function $F$. Hence, $Y_{ni}$ is a random variable whose distribution depends on $F$, while $y_{ni}$ represents its realization. Further, denote $\Gamma_{n}$ as a diagonal $n\times n$ matrix with $i$th entry $\sigma_{ni}=|x_{ni}\theta+z_{ni}\eta|$; this specification allows for relaxation over the homoscedasticity assumption of the units. The only requirement we set is $\eta_{1}\neq 0$.
@@
We introduce some minimal assumptions required for the remainder of the paper. These are not meant to limit practical applicability but to exclude cumbersome or degenerate cases. They are inspired by gutenbrunner1992regression and koenker1999goodness and are standard in inference procedures for quantile regressions.
@@ 3 Closed testing principle
To test the hypotheses of interest $H_{0}^{1},\ldots,H_{0}^{K}$ with FWER control at a pre-specified level $\alpha\in(0,1)$, we employ the closed testing principle (marcus1976closed). This is the optimal framework for controlling the FWER, as any valid multiple testing method is either equivalent to a closed testing procedure or uniformly improved by one. This optimality property extends beyond FWER control and holds for other error measures, such as the false discovery proportion (goeman2021only).
@@
For illustration, consider the following simple example, adapted from goeman2011multiple. Figure 1 shows the tree of all intersection hypotheses for $K=3$, reflecting the subset-superset relationships. Each level corresponds to index sets of the same size, and an arrow from a node $\mathcal{E}$ to a node $\mathcal{C}$ indicates that $\mathcal{C}\subset\mathcal{E}$. The closed testing procedure rejects an individual hypothesis, for example $H_{1}$, only if the local test rejects all intersection hypotheses that include it, namely $H_{\{1\}}=H_{1}$, $H_{\{1,2\}}=H_{1}\cap H_{2}$, $H_{\{1,3\}}=H_{1}\cap H_{3}$, and $H_{\{1,2,3\}}=H_{1}\cap H_{2}\cap H_{3}$.
Figure 1: Tree of intersection hypotheses for three individual hypotheses. Adapted from goeman2011multiple.
In the context of quantile regression, the possible local tests are limited to either tests for individual hypotheses, such as the univariate rank-score test (koenker2005quantile, Section 3.5.4), combined with a Bonferroni correction, or the Wald-type test of koenker2005quantile. However, the Bonferroni correction is known to have notoriously low statistical power (pesarin2001multivariate), while the Wald-type alternative converges slowly to its asymptotic distribution and exhibits distortions in finite samples. These limitations are illustrated, for example, in Section 5 and in de2025closed.
@@ 4 Multivariate rank-score test
The proposed local test relies on a rank-based testing approach. For the hypothesis $H_{0}^{j}$ corresponding to a single quantile of interest, the univariate test based on regression rank-scores is described in koenker2005quantile. This is a score-type test, meaning that the test statistic is derived from the sub-model implied by $H_{0}^{j}$. To our knowledge, the literature has focused exclusively on such univariate hypotheses. This section outlines the theoretical proposal of this manuscript, that is, the extension to intersection hypotheses over $k=1,\dots,K$ quantiles, where the case $k=1$ coincides with the cited univariate test. For simplicity of notation, we will refer to the intersection hypothesis corresponding to the first $k$ quantiles $\{\tau_{1},\ldots,\tau_{k}\}$, that is,
@@
The first key component of the proposed approach is the vector of regression rank-scores, derived from the estimation procedure of the quantile regression model (koenker2005quantile, Section 6.2) and computed under the sub-model implied by the intersection null hypothesis, to which we will refer as the null model. Let $0<\tau<1$ denote a generic quantile level. The regression rank-scores are
@@
As detailed by koenker2005quantile, the score function is obtained by integrating an appropriate score-generating function $\phi:(0,1)\to\mathbb{R}$, which must be non-decreasing and square-integrable. Letting $\hat{a}_{ni}(\tau)$ denote the $i$th component of $\hat{a}_{n}(\tau)$, the score function is thus defined as
@@
While gutenbrunner1993tests introduced general score-generating functions $\phi(\cdot)$ for shift-location problems, koenker1999goodness advocated the use of a quantile-specific function for $\tau$-specific problems:
@@
Let $d_{ni}$ denote the $n$ entries of the vector $D_{n}=X_{n}-\hat{X}_{n}$. Following gutenbrunner1992regression, the quantile rank-score process is
@@
This stochastic process forms the basis of the score-based test statistic proposed below. It is characterized by the following theorem from gutenbrunner1992regression.
@@
and
$T_{n}=S_{n}^{\top}A^{-1}_{n}S_{n},$
where $A_{n}$ is a $k\times k$ matrix whose $(\ell,r)$th entry equals $V_{n}\delta_{\ell r}$ with $\delta_{\ell r}=\,\min\{\tau_{\ell},\tau_{r}\}-\tau_{\ell}\tau_{r}$.
@@
The proof of Theorem 1 in gutenbrunner1992regression shows that
@@
is used to prove Theorem 1 by gutenbrunner1992regression. On the other hand, under the local alternatives, the additional term
@@
For the selection of the bandwidth parameter $h_{n}$, we follow the Hall–Sheather rule (hall1988distribution) as adopted by koenker1999goodness. Since the denominator in (4) may not be positive, $\hat{f}_{ni}(Q_{ni}(\tau)|z_{ni})$ is replaced by
@@
where $\varepsilon>0$ is a small tolerance parameter. This procedure estimates the elements of $\Gamma_{n}$ up to a scale factor independent of $i$, which cancels in the expression for $\hat{X}_{n}$ (koenker1999goodness). The choice of $0.01$ is made for computational reasons and is practically irrelevant.

### v2

@@ Abstract
This paper tackles the challenge of performing multiple quantile regressions across different quantile levels and the associated problem of controlling the familywise error rate, an issue that is generally overlooked in practice. We propose a multivariate extension of the rank-score test and embed it within a closed-testing procedure to efficiently account for multiple testing. Then we further generalize the multivariate test to enhance statistical power against alternatives in selected directions. Theoretical foundations and simulation studies demonstrate that our method effectively controls the familywise error rate while achieving higher power than traditional corrections, such as Bonferroni.
@@ 1 Introduction
We focus on the problem of testing the evidence of a covariate effect across multiple quantile levels. In the literature, this is often referred to as regional quantile testing when the quantile levels form a continuous range, and as global or partial quantile regression when the levels form a finite set of quantiles $\{\tau_{1},\dots,\tau_{K}\}$. Various methods exist for regional quantile testing, with examples found in the work of Sun and He (2021); Belloni et al. (2019); He et al. (2010). In contrast, a valid approach for conducting quantile regressions at multiple discrete $K$ levels is lacking. Typically, researchers perform $K$ separate regressions, analyzing each regression’s results and the corresponding $p$-values/test statistics independently.
@@
This practice, however, leads to a multiple testing problem (Goeman and Solari, 2014), where performing $K$ separate tests increases the risk of false discoveries if $p$-values are not adjusted to control family errors such as familywise error (FWER) or false discovery rates. The multiplicity problem is not commonly addressed within the quantile regression literature. To the best of our knowledge, only Mrkvicka et al. (2023) consider multiple testing in this framework; yet, their proposal lacks a solid theoretical foundation.
@@
The strategy of applying quantile regression across various quantiles and making inferences about the effects of covariates without addressing multiple testing issues is a common oversight in various disciplines. Numerous examples can be found in studies spanning markets (Wang et al., 2012), biology (Agarwal et al., 2019), management (Bien et al., 2016; Karimi Alavijeh and Zangoei, 2024), economics (Fousekis and Lazaridis, 2005; Falk, 2012), public health (Juvanhol et al., 2016), health (Cheah et al., 2022), education (Bassett Jr et al., 2002), energy (Cheng et al., 2020), and climate (Chamaillé-Jammes et al., 2007). This widespread practice highlights a methodological gap, emphasizing the need for the field to develop and apply rigorous statistical methods that properly account for multiple comparisons and ensure the validity of the findings (Ioannidis, 2005).
@@
We propose a principled yet straightforward framework for performing valid inference across multiple quantile levels. A naive but valid approach would be to apply standard corrections, such as Bonferroni, to the set of $p$-values obtained from the $K$ separate quantile regressions. However, as demonstrated by Goeman et al. (2021), only closed testing procedures are admissible for strong FWER control; every other method is either equivalent to, or uniformly dominated by, some closed testing procedure. Consequently, we adopt the closed testing principle as the inferential backbone of our approach. The closed testing framework controls the FWER by testing all possible intersections of the individual hypotheses $H_{0}^{1},\dots,H_{0}^{K}$ using valid local tests. In short, for any $j=1,\dots,K$, we can reject $H_{0}^{j}$ at a pre-determined level $\alpha$ if all intersection hypotheses containing $H_{0}^{j}$ are rejected at that level $\alpha$. The validity and efficiency of the closed testing procedure thus hinge on the choice of these local tests. The commonly used Wald-type test for quantile regression (Koenker and Portnoy, 1996) suffers from finite-sample distortions, especially when heteroscedasticity affects the data, and slow convergence to its asymptotic distribution, even under independent and identically distributed errors (De Santis et al., 2025; Koenker, 2005).
@@
A robust alternative is the rank-score test (Gutenbrunner et al., 1993; Koenker and Machado, 1999; Kocherginsky et al., 2005), whose multivariate version, however, has not yet been fully developed and theoretically investigated. We therefore first propose the multivariate extension of the rank-score test and study its theoretical proprieties. We then develop a generalization of the multivariate rank-score test, based on alternative weighting matrices, and integrate the resulting class of statistics as the local test within the closed testing procedure. This approach increases analytical flexibility and guarantees valid control of type I errors across multiple quantile levels, thereby improving the reliability of statistical inferences from quantile regression analyses.
@@
To facilitate the application of the proposed methodology, we provide an R package named quasar (QUAntile Selective inference Adjustment via Rank-scores) available on CRAN at https://CRAN.R-project.org/package=quasar.
@@
The manuscript is organized as follows. Section 2 presents the necessary background on quantile regression, while Section 3 introduces the closed testing procedure. Section 4 details the theory of the multivariate version of the rank-score test. Section 5 extends the multivariate rank-score test by introducing a class of generalized test statistics based on alternative weighting matrices. Section 6 presents simulation studies showing that the Wald-type test fails to control the type I error rate, whereas the proposed method effectively controls the FWER while maintaining competitive power compared to traditional corrections such as Bonferroni.
@@ 2 Setting
We assume that the following location-shift model (Koenker and Machado, 1999) generates the data:
@@
where $\beta,\theta\in\mathbb{R}$, $\gamma,\eta\in\mathbb{R}^{p}$, and $U_{ni}$ are independent and identically distributed zero-mean errors from a distribution function $F$. Hence, $Y_{ni}$ is a random variable whose distribution depends on $F$, while $y_{ni}$ represents its realization. Further, denote $\Gamma_{n}$ as a diagonal $n\times n$ matrix with $i$th entry $\sigma_{ni}=|x_{ni}\theta+z_{ni}\eta|$; this specification allows for relaxation over the homoscedasticity assumption of the units. The only requirement we set is $\eta_{1}\neq 0$.
@@
We introduce some minimal assumptions required for the remainder of the paper. These are not meant to limit practical applicability but to exclude cumbersome or degenerate cases. They are inspired by Gutenbrunner and Jureckova (1992) and Koenker and Machado (1999) and are standard in inference procedures for quantile regressions.
@@ 3 Closed testing principle
To test the hypotheses of interest $H_{0}^{1},\ldots,H_{0}^{K}$ with FWER control at a pre-specified level $\alpha\in(0,1)$, we employ the closed testing principle (Marcus et al., 1976). This is the optimal framework for controlling the FWER, as any valid multiple testing method is either equivalent to a closed testing procedure or uniformly improved by one. This optimality property extends beyond FWER control and holds for other error measures, such as the false discovery proportion (Goeman et al., 2021).
@@
For illustration, consider the following simple example, adapted from Goeman and Solari (2011). Figure 1 shows the tree of all intersection hypotheses for $K=3$, reflecting the subset-superset relationships. Each level corresponds to index sets of the same size, and an arrow from a node $\mathcal{E}$ to a node $\mathcal{C}$ indicates that $\mathcal{C}\subset\mathcal{E}$. The closed testing procedure rejects an individual hypothesis, for example $H_{1}$, only if the local test rejects all intersection hypotheses that include it, namely $H_{\{1\}}=H_{1}$, $H_{\{1,2\}}=H_{1}\cap H_{2}$, $H_{\{1,3\}}=H_{1}\cap H_{3}$, and $H_{\{1,2,3\}}=H_{1}\cap H_{2}\cap H_{3}$.
Figure 1: Tree of intersection hypotheses for three individual hypotheses. Adapted from Goeman and Solari (2011).
In the context of quantile regression, the possible local tests are limited to either tests for individual hypotheses, such as the univariate rank-score test (Koenker, 2005, Section 3.5.4), combined with a Bonferroni correction, or the Wald-type test of Koenker (2005, Section 3.5.3). However, the Bonferroni correction is known to have notoriously low statistical power (Pesarin, 2001), while the Wald-type alternative converges slowly to its asymptotic distribution and exhibits distortions in finite samples. These limitations are illustrated, for example, in Section 6 and in De Santis et al. (2025).
@@ 4 Multivariate rank-score test
The proposed local test relies on a rank-based testing approach. For the hypothesis $H_{0}^{j}$ corresponding to a single quantile of interest, the univariate test based on regression rank-scores is described in Koenker (2005, Section 3.5.4). This is a score-type test, meaning that the test statistic is derived from the sub-model implied by $H_{0}^{j}$. To our knowledge, the literature has focused exclusively on such univariate hypotheses. This section outlines the theoretical proposal of this manuscript, that is, the extension to intersection hypotheses over $k=1,\dots,K$ quantiles, where the case $k=1$ coincides with the cited univariate test. For simplicity of notation, we will refer to the intersection hypothesis corresponding to the first $k$ quantiles $\{\tau_{1},\ldots,\tau_{k}\}$, that is,
@@
The first key component of the proposed approach is the vector of regression rank-scores, derived from the estimation procedure of the quantile regression model (Koenker, 2005, Section 6.2) and computed under the sub-model implied by the intersection null hypothesis, to which we will refer as the null model. Let $0<\tau<1$ denote a generic quantile level. The regression rank-scores are
@@
As detailed by Koenker (2005), the score function is obtained by integrating an appropriate score-generating function $\phi:(0,1)\to\mathbb{R}$, which must be non-decreasing and square-integrable. Letting $\hat{a}_{ni}(\tau)$ denote the $i$th component of $\hat{a}_{n}(\tau)$, the score function is thus defined as
@@
While Gutenbrunner et al. (1993) introduced general score-generating functions $\phi(\cdot)$ for shift-location problems, Koenker and Machado (1999) advocated the use of a quantile-specific function for $\tau$-specific problems:
@@
Let $d_{ni}$ denote the $i$th entry of the vector $D_{n}=X_{n}-\hat{X}_{n}$. Following Gutenbrunner and Jureckova (1992), the quantile rank-score process is
@@
This stochastic process forms the basis of the score-based test statistic proposed below. It is characterized by the following theorem from Gutenbrunner and Jureckova (1992, Theorem 1(iii)).
@@
and the $k\times k$ matrix $\Delta$ whose $(\ell,r)$th entry is $\delta_{\ell r}=\,\min\{\tau_{\ell},\tau_{r}\}-\tau_{\ell}\tau_{r}$. The test statistic for the local test is
$T_{n}=S_{n}^{\top}A^{-1}_{n}S_{n},$ (2)
where $A_{n}=V_{n}\Delta$, and $V_{n}$ is the scalar introduced in Assumption 7.
@@
The proof of Theorem 1 in Gutenbrunner and Jureckova (1992) shows that
@@
is used to prove Theorem 1 by Gutenbrunner and Jureckova (1992). On the other hand, under the local alternatives, the additional term
@@
For the selection of the bandwidth parameter $h_{n}$, we follow the Hall–Sheather rule (Hall and Sheather, 1988) as adopted by Koenker and Machado (1999). Since the denominator in (5) may not be positive, $\hat{f}_{ni}(Q_{ni}(\tau)|z_{ni})$ is replaced by
@@
where $\varepsilon>0$ is a small tolerance parameter. This procedure estimates the elements of $\Gamma_{n}$ up to a scale factor independent of $i$, which cancels in the expression for $\hat{X}_{n}$ (Koenker and Machado, 1999). The choice of $0.01$ is made for computational reasons and is practically irrelevant.
5 Generalized multivariate rank-score test
While the test described in the previous section is appealing for its simple mathematical formulation and reference distribution, its power can be significantly improved without compromising the type I error control.
To begin with, consider the distribution of the test statistic under alternatives, in the setting of Theorem 3. Once the noncentrality parameter is made explicit, the limitations in terms of statistical power become evident. Consider, for example, the case $k=2$; the noncentrality parameter is
$\zeta=g_{1}^{2}a_{11}+2g_{1}g_{2}a_{12}+g^{2}_{2}a_{22}$
where $A=V\Delta$ and $a_{\ell r}$ denotes the $(\ell,r)$th entry of the matrix $A^{-1}$. Thus, the power of the test decreases as a function of the correlation among the test statistics. Furthermore, the finite-sample counterpart of $A$ is $A_{n}=V_{n}\Delta$, where $\Delta$ is deterministic and depends only on the chosen quantile levels.
It is possible to replace the matrix $A_{n}^{-1}$ in the definition of the test statistic (2) with an alternative weighting matrix. The theory of multivariate score tests with different weighting matrices is developed in Goeman et al. (2006, 2011) for mean regression models. Motivated by this framework, we introduce an alternative test statistic
$T_{n}^{*}=S_{n}^{\top}BS_{n},$
where $B$ is any symmetric, positive-definite, $k\times k$ matrix. Under the null and alternative hypotheses, the original statistic $T_{n}$ converges in distribution to a (noncentral) $\chi^{2}_{k}$ random variable (Theorems 2-3). By contrast, $T_{n}^{*}$ converges to a weighted sum of $k$ independent (noncentral) $\chi^{2}_{1}$ random variables:
Theorem 4.
Let Assumptions 17 hold and suppose that $H_{0}^{\{1,\ldots,k\}}$ is true. Then $T_{n}^{*}\xrightarrow{d}T^{*}\sim\sum_{i=1}^{k}\lambda_{i}\chi^{2}_{1}$, where the weights $\lambda_{i}$ are the eigenvalues of the product matrix $BA$.
Proof.
From Theorem 2, $S_{n}\xrightarrow{d}S\sim N(0,A)$. Hence, we may write $A^{-1/2}S=S_{z}\sim N(0,I_{k})$. Then
$T^{*}=S^{\top}BS=S^{\top}_{z}\left(A^{1/2}BA^{1/2}\right)S_{z}=\sum_{i=1}^{k}\lambda_{i}(b_{i}^{\top}S_{z})^{2},$
where $b_{i}$ denote the eigenvectors of the matrix $A^{1/2}BA^{1/2}$. Notice that, by the rotational invariance of the multivariate normal distribution, and being the eigenvectors orthogonal by definition, the random vector $(b_{1}^{\top}S_{z},\ldots,b_{k}^{\top}S_{z})^{\top}$ has components which are independent standard normal random variables.
The final claim follows from the fact that, given the assumptions made, the eigenvalues of $A^{1/2}BA^{1/2}$ and $BA$ coincide. ∎
The reference distribution cannot be derived analytically, but it can be approximated numerically, for instance using the method proposed by Imhof (1961). Further, by the definition of $B$, the weights $\lambda_{i}$ of the distribution are all strictly positive.
While the choice of $B$ does not affect the control of the type I error, it is crucial for the power properties of the test. When testing multiple parameters simultaneously, it is generally not possible to maximize power against all alternatives (Goeman et al., 2006). Indeed, if the region of alternatives is viewed as a $k$-dimensional ball centered at the null, the intersection test is primarily sensitive along a particular direction. The choice of $B$ effectively rotates this direction, influencing which alternatives the test is most powerful against. The following theorem derives the distribution of the test statistic under local alternatives.
Theorem 5.
Let Assumptions 17 hold and suppose that $H_{n}^{\{1,\ldots,k\}}$ is true. Then $T_{n}^{*}\xrightarrow{d}T^{*}\sim\sum_{i=1}^{k}\lambda_{i}\chi^{2}_{1}(\zeta_{i})$, where $\zeta_{i}=(b_{i}^{\top}g_{z})^{2}$ and $g_{z}=A^{-1/2}g$.
Proof.
The proof proceeds along the same lines as that of Theorem 4. From Theorem 3, $S_{n}\xrightarrow{d}S\sim N(g,A)$. Then
$T^{*}=(S_{z}+g_{z})^{\top}A^{1/2}BA^{1/2}(S_{z}+g_{z})=\sum_{i=1}^{k}\lambda_{i}[b_{i}^{\top}(S_{z}+g_{z})]^{2}$
where the elements $\lambda_{i}$ and $b_{i}$ are, respectively, the eigenvalues and eigenvectors of the matrix $A^{1/2}BA^{1/2}$. The properties of the normal distribution apply as in Theorem 3, since each element $b_{i}^{\top}g_{z}$ applies as a location shift of a standard normal random variable. ∎
We conclude this section with some practical remarks. The choice of the weighting matrix clearly affects the distribution of the test statistic under alternatives in multiple ways, influencing both the noncentrality parameters and the weights of the chi-square-based reference distribution. Moreover, the convergence rate to the asymptotic distribution depends on the selected quantile levels. Hence, defining an optimal weighting matrix is not straightforward, even when prioritizing specific quantiles, while simultaneously accounting for the multiplicity inherent in the global testing problem.
Consider the identity matrix $I_{k}$ as the weighting matrix, and $S$ to be a generic zero-mean normal random vector with positive-definite variance. Then $(S-g)^{\top}(S-g)$ is a weighted Euclidean distance, with weights given by the variance matrix of $S$. By performing a singular value decomposition of this matrix,
$(S-g)^{\top}(S-g)=\sum_{i=1}^{k}\lambda_{i}\left(S_{i}+\frac{b_{i}^{\top}g}{\sqrt{\lambda_{i}}}\right)^{2}.$
If the variance matrix is equicorrelated with identical diagonal elements, the power is equally distributed across the elements of $g$. In this scenario, the overall power depends on the magnitude of the correlation, as can be seen from the explicit form of the eigenvalues and eigenvectors. In this case, the first eigenvector has identical entries, the sum of the eigenvalues is constant, and the first eigenvalue increases with the correlation. If this eigenvalue dominates the others, the distance induced by $g$ is mainly in the first principal component. The test is primarily sensitive in this direction, leading to increased power. Heterogeneity in either the diagonal or off-diagonal elements of the variance matrix influences both the distribution of power among the components and the overall average power.
Returning to the quantile regression framework, the variance matrix depends only on the quantile levels, up to the scalar factor, and is therefore a deterministic, known quantity. Hence, differences in variances across components affect the distribution of the power among the elements of $g$, penalizing the extreme quantiles. To mitigate this issue, different weighting matrices can be used. This should be viewed as distributing the overall power differently, rather than providing a possible maximization over the entire set of quantiles. For example, to correct for the imbalance induced by heterogeneous variances, one may use the weighting matrix $\mbox{diag}\{\Delta\}^{-1}$. Another natural choice arises when assumptions are made on the distribution of the errors. In this case, the elements of the vector $g$ depend on the error density computed at the target quantiles. To balance power across the alternatives $h(\tau_{j})$, the diagonal of the weighting matrix can be set equal to the reciprocals of these densities. However, if the assumed error distribution is misspecified, this procedure does not properly redistribute power across the true alternatives, but rather across a more complex combination of the true (unknown) distribution and the assumed one.